OpenStudy (anonymous):

find the average value of y=(x^2)[sq root of (x^3+1)]

OpenStudy (anonymous):

PLEASE HELP ME! IM SO FRUSTRATED :((

OpenStudy (nikvist):

what is interval?

OpenStudy (anonymous):

[0,2]

OpenStudy (nikvist):

\[\frac{1}{2}\int\limits_0^2x^2\sqrt{x^3+1}dx=\quad;\quad y=x^3+1,dy=3x^2dx\] \[=\frac{1}{2}\int\limits_1^9\sqrt{y}\frac{dy}{3}=\frac{1}{6}\cdot\frac{2}{3}y^{3/2}|_{1}^{9}=\frac{1}{9}(27-1)=\frac{26}{9}\]

OpenStudy (anonymous):

thank you so much! can you help me on one more with average value please? a function f(t)=6+cos(t/10)=3sin(7t/40), is used to model the velocity of a plane in miles per minute. according to this model, what is the average velocity of the plane for 0 less than or equal to "t" less than or equal to 40?